OpenStudy (anonymous):

Solve for x & show all steps: 2^x = 64 5^2x = 5^10 Now, I obviously know the answers to these but my teacher marks hard as hell so I need to know the steps behind it because she assumes I'm guessing.

OpenStudy (anonymous):

i dont think theres really steps behind the first one ;-;

OpenStudy (anonymous):

she's a horrid marker and im sure there aren't even steps behind it

OpenStudy (anonymous):

For the Second, Since they have an equivalent base of 5, you just solve the exponents. so cancel out the 5's and solve 2x=10. divide by 2 and get x=10.

OpenStudy (anonymous):

x=5* o-o Omg.

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

ok, got it. i'll apply this to the test tomorrow and if she marks it like she has been im just gonna confront her lol

OpenStudy (anonymous):

Okay xD GOODLUCK c:

OpenStudy (phi):

make both sides have the *same* base then you can equate the exponents

OpenStudy (anonymous):

ok so what if the bases are different like what if i had 5^3x = 7^10 or something, how would i go about doing that

OpenStudy (anonymous):

sorry about all these random questions, school just started and it's not fresh in my mind yet

OpenStudy (phi):

for the 2nd problem, you already are set \[ 5^{2x} = 5^{10} \\ 2x= 10 \\ x= 5 \]

OpenStudy (phi):

if you have different bases, then you use a more "sophisticated" idea. take the log of each side. \[ 5^{3x} = 7^{10} \\ \log\left(5^{3x} \right)= \log\left(7^{10} \right) \\ 3x \log(5) = 10 \log(7) \\ x = \frac{ 10 \log(7)}{3 \log(5)} \]

OpenStudy (phi):

for the first problem 2^x = 64 notice 64 is a power of 2. (If you did not know, you could hope it's true, and divide by 2 to get 2*32, and then 2*2*16, and then 2*2*2*8, and then 2*2*2*2*2*2 = 2^6

OpenStudy (anonymous):

oh okay, that seems good to do

OpenStudy (phi):

it might be slightly trickier if you see \[ 9^x = 27^2 \] 27 is not a power of 9. the next power 9^2 = 9*9=81 *but*, notice we could factor 9 into 3*3= 3^2 and 27 is 3*3*3 = 3^3 and we have \[ \left(3^2\right)^x = \left( 3^3\right)^2 \] use this property (good to know) \[ \left(a^b\right)^c = a^{bc} \] to write \[ 3^{2x}= 3^{3 \cdot 2}\\ 3^{2x}= 3^6\] and 2x=6 , and x= 6/2 = 3

OpenStudy (triciaal):

2^x = 64